factor the quadratic expression completely, and find the roots of the expression.
1. 135x2 - 222x + 91
2. 6x2 - 42x + 72

Answer by checkley75(3666)

135X^2-222X+91=0
(9X-7)(15X-13)=0
9X-7=0
9X=7
X=7/9 ANSWER.
15X-13=0
15X=13
X=13/15 ANSWER.
--------------------------------------------------
6X^2-42X+72=0
(3X-9)(2X-8)=0
3X-9=0
3X=9
X=3 ANSWER.
2X-8=0
2X=8
X=8/2
X=4 ANSWER.